PHOTOION PRODUCTION OF HFH^{+} FROM (HF)_{2}

\underline{QICHI HU}, JOHN W. HEPBURN, Department of Chemistry, University of British Columbia, Vancouver, BC V6T 1Z1, Canada.

Total photoion yield and threshold production of HFH^{+} from (HF)_{2} has been studied by single photon VUV excitation. The VUV was scanned in the energy range from 15.9 to 14.6 eV. In this energy range there are two allowed channels producing HFH^{+} ion: (HF)_{2} + hν → HF H^{+} + e^{−} + F and (HF)_{2} + hν → HF H^{+} + F^{−}. The first channel was found to be the dominant one. Vibrational bands of HF H^{+} ion (ν^{+} = ν_{1} + ν_{2} + ν_{3} up to 6) were assigned in the threshold photoionization spectrum, while clearly resolved sub-vibrational structures were also observed. The threshold of (HF)_{2}(ν′′ = 0) + hν → HF H^{+}(ν^{+} = 0) + e^{−} + F was determined to be 14.72 ± 0.02 eV. Our data shows that the HF H^{+} ion production is greatly enhanced by Rydberg resonances of (HF)_{2}^{∗}, which act as the intermediate states in the production of HF H^{+} ions.